OpenStudy (anonymous):

Help help help!!! Can someone help me and explain step by step how to multiply this:

OpenStudy (anonymous):

\[\frac{ x ^{3}+1 }{ x ^{3}-x ^{2}+x }\times \frac{ 6x }{ -54x-54 }\]

OpenStudy (valpey):

Do you agree that:?\[\frac{AB}{CD}*\frac{EF}{GH}=\frac{ABEF}{CDGH}\]

OpenStudy (valpey):

And do you agree that:?\[x^3+1=(x+1)(x^2-x+1)\]\[x^3−x^2+x=x(x^2-x+1)\]\[-54x-54=-54(x+1)\]\[6x=6*x\]

OpenStudy (anonymous):

Yes, I do. Sorry, I was away. @Valpey

OpenStudy (anonymous):

@Hero I only have 3 more questions. ): Please help?

hero (hero):

Did you get that previous question correct?

OpenStudy (anonymous):

No, because I didn't know how to figure out what it equaling with the square root.

hero (hero):

Let's go back to it shall we?

OpenStudy (anonymous):

Alright, well, I mean, we left off at \[\sqrt{0}=\sqrt{(x-5)^{2}}\]

hero (hero):

Yes, and what is the square root of zero?

hero (hero):

It's okay to say you don't know

hero (hero):

Okay, how about this: \(0^2 = 0\) right?

OpenStudy (anonymous):

Yes.

hero (hero):

Okay, so we can rewrite the equation like this: \[\sqrt{0^2}=\sqrt{(x-5)^{2}}\] And guess what? The square root and the square are inverses of each other, so they cancel. So what are you left with?

OpenStudy (anonymous):

0=x-5

hero (hero):

Precisely. So continue solving for x.

OpenStudy (anonymous):

Thank you!! Okay, so another question is the one al the way at the top, want me to retype the equation?

hero (hero):

I could have sworn I've seen that equation before.

OpenStudy (anonymous):

You have but we never worked it out. Oh! And what exactly is a domain?

hero (hero):

The domain of a function is the set of all x values that are valid for the function.

hero (hero):

We should probably go back to that question. I could have sworn I posted something to help you.

hero (hero):

I suppose you don't know how to get back to your own question.

OpenStudy (anonymous):

I factored everything, but from there, I have no clue.

hero (hero):

I'm going back to the other question. Are you coming with me?

OpenStudy (anonymous):

The one at the top?

hero (hero):

I'll just post a link to it.

OpenStudy (anonymous):

Okay